Project

General

Profile

20170502 » History » Revision 8

Revision 7 (Boyd, Steven, 02 May 2017 12:55) → Revision 8/12 (Boyd, Steven, 02 May 2017 12:55)

h1. Ops/Comm Scheduling meeting : 2/5/2017 

 Location : CR21 at 14:00 pm 

 Phone access :    *New* phonebridge number 
 UK: 08082380274 
 US: 631.621.5253 
 Access code: 1779339517 

 h2. Agenda 

 * Actions 
 ** Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands 
 ** Macwaters - Re-make dodgy compressor2 remote cable connector - stands 
 ** Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress 
 ** Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress 
 ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress 
 ** Macwaters - Report on status of FC psu repair (stands) 
 *** *This needs to be cross-checked* - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it _in situ_ before the run? 
 *** _Josef :"we just run 50A through it when it is connected to the magnet"_  
 ** Hall Manager - water system change-over procedure (stands) 
 ** EOverton - Status of @miceraid5@ disk cleanup (stands) 
 ** JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands) 
 ** CWhyte to contact DL to clarify their [[Hall_Task_List_Spring17|Hall Task List]] items 
 ** The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not 
 *** _Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC_ 
 ** HNebrensky to send BLOC-chasing email 

 * DC Update 
 * LH2 Update  
 * Tracker Update 
 * [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin)  
 * Schedule for startup 
 ** Startup run on Monday May 8th - Hall should be reserved for operations activities all day. Controlled Entry is possible 
 but only when permitted by the MOM. 
 ** We will *not* be running the SS magnets this user cycle, or taking data with Neon. 
 ** We will need to run the FC at some point. 
 ** Co-ordination between DC and MOM will be key. 

 AOB: 

 * Last meeting : [[20170425|Tuesday Apr 25 2017]] 
 * Next meeting : [[20170507|Tuesday May 7 2017]]